# LightGaussian: Unbounded 3D Gaussian Compression with 15x Reduction and 200+ FPS

### v4

@@ 3.3 Distilling into Compact SHs
In the uncompressed Gaussian Splat data, a significant portion is comprised of Spherical Harmonics (SH) coefficients, requiring (45+3) floating-point values per splat, which represent 81.3 percent of the total attribute volume. Reducing the degree of SH, while beneficial for decreasing disk space usage, leads to a noticeable loss of surface ‘shininess’, particularly affecting specular reflection variance when the viewpoint is altered.
To achieve a balance between model size and scene reflectance quality, we propose transferring knowledge from well-trained high-degree SHs to their compact counterparts (lower-degree) via distillation. Specifically, we distill the teacher model (with full-degree SHs, and is frozen) to the student model (with truncated lower-degree SHs). The supervisions are from the difference of the predicted pixel intensities $\boldsymbol{C}$ from the two models :
@@
Synthesize Pseudo Views
The specular reflection rays are reflected off the surface of the light source when we move the viewpoints. We can augment the training views by sampling additional viewpoints that can reflect such reflectance. As we know the camera locations and viewing directions of all training views, we randomly sample pseudo views around each of the training views, obeying a Gaussian distribution, while fixing the camera view directions:
@@
3.4 VecTree Attribute Compression
Vector quantization is well-explored in voxel-based NeRF representations [38, 66, 56, 70]. Although a high compression rate can be reached, Gaussian attributes cannot be easily quantized in the same fashion. The reason is that each attribute of the explicit representation has its own physical meaning and complex correlation (e.g., the rotation and scaling formulate the covariance matrix), and we empirically find that certain attributes such as opacity, rotation, and scale are more sensitive to such discretized representation, with a dramatic drop in accuracy. Therefore, we propose two ways to mitigate this gap: first, we reutilize the significance score from Sec. 3.2 and perform excessive vector quantization on the SHs of the least important Gaussians. For the Gaussian location, we draw from point cloud compression and adopt octree-based lossless compression into our framework. For the remaining important SHs and other attributes representing the Gaussian shape, rotation, and opacity, we save these features in float16 format.
Vector Quantization
VQ partitions the Gaussians $\mathcal{G}=\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{N}\}$ to the K codes in the codebook $\mathcal{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{K}\}$, where each $\mathbf{g}_{j},\mathbf{c}_{k}\in\mathbb{R}^{d}$ and K $\ll$ N. $d$ means attributes’ dimension. Specifically, $\mathcal{C}$ is initialized via K-means. Then, VQ iteratively samples a batch of $\mathcal{G}$, associates them to the closest codes by euclidean distance, and updates each $\mathbf{c}_{k}$ via moving average rule: $\mathbf{c}_{k}=0.8\cdot\mathbf{c}_{k}+(1-0.8)\cdot 1/\mathbf{T}_{k}\cdot\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}\mathbf{GS}_{j}\cdot\mathbf{g}_{j}$, where $\mathbf{T}_{k}=\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}{\mathbf{GS}_{j}}$ is the total significance score assigned to the code vector $\mathbf{c}_{k}$, $\mathcal{R}(\mathbf{c}_{k})$ is the set of Gaussians associated to the $k$-th code.
@@ 3.4 VecTree Attribute Compression
Plenoxels [67] 6.79 2100MB 23.62 0.670 0.443 11.2 2700MB 21.07 0.719 0.379
INGP-Big [45] 9.43 48MB 26.75 0.751 0.299 2.79 48MB 21.92 0.744 0.305
Mip-NeRF 360 [5] 0.06 8.6MB 29.23 0.844 0.207 0.09 8.6MB 22.21 0.759 0.257
VQ-DVGO [38] 6.53 56.07MB 25.44 0.699 0.325 - - - - -
3D-GS [32] 139 727MB 29.13 0.870 0.185 106 380MB 23.11 0.822 0.219
Ours 215 42.48MB 28.45 0.857 0.210 209 22.43MB 22.83 0.807 0.242
Table 1: Quantitative Comparisons in Real-world Large-scale Scenes. We compare LightGaussian with the original 3D-GS [32], efficient voxel-based NeRFs [67, 45], voxel NeRF with vector quantization [38], and compact MLP-based NeRF [5]. Voxel-based methods all exhibit a lack of sufficient capacity for representing large-scale scenes and are not able to run at real-time speed. Mip-NeRF 360 produces the best visual quality, but it requires more than 16s to render a single image. Our method achieves a good balance among FPS, model size, and rendering quality.
Figure 4: Visual Comparisons. We compare LightGaussian with the vanilla 3D-GS [32], presenting a residual map between prediction and ground-truth scaled from 0 to 127 to highlight the differences. We observe that LightGaussian preserves the specular reflections (yellow boxes) after converting to a compact format. Additionally, a slight lightness change is noted after the conversion, as shown in the bottom white box. For dynamic viewpoint comparisons, please refer to our supplementary video material.
Octree-based Compression
To deal with the unstructured and high-precision 3D points location, we avoid the lossy compression which is sensitive to the subsequent rasterization accuracy. Drawing inspiration from the Point Cloud Compression literature [52], which applies lossless encoding on the Gaussian location, utilizing the octree geometry codec in G-PCC [44] consumes only a very small amount of bits.
@@ 4.1 Experimental Settings
Datasets and Metrics
Comparisons are performed on the widely adopted scene-scale view synthesis data created by Mip-NeRF360 [6], which contains nine real-world large-scale scenes (5 unbounded outdoors, 4 indoors with detailed backgrounds). We utilize the published seven scenes for comparison following the train/test split in Mip-NeRF360 [6]. Additionally, we use another large-scale unbounded dataset, the Tanks and Temples dataset [34], where we perform comparisons using the same two scenes as used in 3D-GS [32]. In total, the adopted datasets consist of scenes with very different capture styles and cover both bounded indoor scenes with detailed backgrounds and large unbounded outdoor cases. The rendering quality is reported using metrics such as peak signal-to-noise ratio (PSNR), structural similarity (SSIM), and perceptual similarity via LPIPS [69].
Compared Baselines
We compare with methods that can model large-scale scenes, including Plenoxel [67], Mip-NeRF360 [5], and 3D-GS [32], along with representative efficient methods such as Instant-ngp [45], which uses a hash grid for storage, and VQ-DVGO [38], utilizing DVGO [55] as the basic representation and introducing voxel pruning and vector quantization to it.
@@
[7]  + Codebook Quant. 300.60 23MB 31.26 0.917 0.220
[8]  + Octree Compression. 300.60 20MB 31.26 0.917 0.220
[9] LightGaussian (Ours) 300.60 20MB 31.26 0.917 0.220
Table 2: Ablation studies on the Gaussian Pruning, SH Compactness, and the VecTree Compression. Scene: Room. Zero-shot Gaussian pruning leads to a degradation in rendering quality (#2), but Co-adaptation can recover most of the scene details (#3). Directly eliminating high-order SH negatively affects the quality (#4), while distillation with pseudo-view helps to mitigate the gap (#5, #6). Codebook quantization significantly reduces the required model size (#7), while lossless Octree compression helps to preserve the overall quality.
@@
Table 3: Ablation study of the Gaussian Pruning, by using different Gaussian attributes for computing its global significance score. By considering only the hit count of each Gaussian from training rays, the zero-shot pruning leads to inferior performance. Incorporating the opacity and volume drives us to a better criterion. The subsequent Gaussian Co-adaptation is used to recover most of the information loss from the pruning of redundant Gaussians.
@@
+FP16 36.76MB 31.58 0.923 0.212
 + VQ All att. 18.58MB 22.97 0.750 0.378
 + VQ All att. $\times$ GS 18.58MB 26.39 0.830 0.327
 + VQ SH. 22.53MB 30.94 0.910 0.225
 + VQ SH $\times$ GS 22.53MB 31.26 0.917 0.220
VecTree 20.65MB 31.26 0.917 0.220
Table 4: Ablation study of hybrid VecTree Quantization (VQ). By quantizing all attributes to FP16, except for Gaussian locations due to their sensitivity, a smaller model is achieved. VQ is applied to all attributes leads to inferior modeling accuracy, but this can be mitigated by using Global Significance (GS) on the least crucial Gaussians. Other attributes (e.g., scale) are also sensitive to VQ, hence we only apply VQ on the SH. Combining this with Octree compression results in our VecTree Compression (the last row), which demonstrates a good balance between size and quality.
Implementation Details
The framework is developed using Pytorch and reuses the differential Gaussian rasterization presented by 3D-GS [32]. We use our own trained checkpoints that match performance with those detailed in the 3D-GS paper [32]. For the Global Significance Calculation, we set the power to 0.1 in Eq. 5, and fine-tune for 5,000 steps in Gaussian Co-adaptation. For the SHs Distillation, we reduce the 3-degree SHs to 2-degree, eliminating 21 elements for each Gaussian, and enhance it with $\sigma$ set to 0.1 in the pseudo view synthesis. For the VecTree Compression, we set the codebook size to 8192.
Figure 5: Visual Comparisons for Ablation Study. We visualize the rendered RGB images and the residual map between the ground-truth image, aligned with the experiment ID as shown in Tab. 2. The final model (Exp #7, #8, #9) demonstrates close results to 3D-GS (Exp #1), while the Gaussian Co-adaptation, along with SH distillation, almost completely mitigates the information loss.
@@ 4.2 Experimental Results
Quantitative Results
To assess the quality and model size of different methods for novel view synthesis, we summarize the quantitative numbers in Tab. 1, including highly efficient voxel-based NeRFs (Plenoxel[67], Instant-NGP [45]), the compact MLP-based Mip-NeRF360 [5], NeRF with vector quantization [38], and 3D Gaussian Splatting [32]. Specifically, in the Mip-NeRF360 datasets, the NeRF-based methods, benefiting from their compact representation by using MLPs, show competitive accuracy, but the slow inference speed (0.06FPS) makes them impractical for real-world applications. Voxel-based NeRFs partly resolve the rendering efficiency issue, but the FPS still lags behind for practical usage, while Plenoxel requires 2.1GB for storing a large-scale scene. VQ-DVGO optimizes the storage issue, but their pruning and purely vector quantization make it challenging to generalize to complex large-scale scenes. On the other hand, 3D-GS strikes a balance in rendering quality and real-time rendering speed; however, it requires nearly gigabytes for storing a single scene. Our pipeline, LightGaussian, achieves the fastest rendering speed (1.55$\times$) compared to all existing methods, thanks to its pruning of insignificant Gaussians for an efficient rasterization process. Moreover, the compact SH representation and VecTree compression further reduce the model redundancy in 3D Gaussians, shrinking the model size from 727MB to 42.48MB on Mip-NeRF360 datasets, a 17$\times$ reduction ratio. LightGaussian also achieves nearly 2$\times$ faster rendering speed on the Tank & Temple datasets, while reducing the storage from 380MB to 22.43MB.
Qualitative Results
We compared the rendering results between 3D-GS and ours, with an emphasis on intricate details and background regions as illustrated in Fig. 4, to highlight the differences. It is observable that 3D-GS and our LightGaussian achieve similar visual quality, even in thin structures within the scenes (the first and third rows). The Gaussian compression method slightly affects the specular effect on the surface (see the second and the fourth rows).
@@ 4.3 Ablation Studies
We ablate components of our method by separately analyzing the proposed modules. Specifically, we find that Iterative Gaussian Pruning is effective in removing redundant Gaussians, SHs Distillation effectively reduces the SH’s degree, and Vector Quantization efficiently compresses the feature space, leading to a more compact and efficient representation. These individual contributions collectively enhance the overall performance of our framework.
Overall Analysis
We report the performance of the investigation on the proposed modules in Tab. 2 and Fig. 5. We verify that our design of Gaussian Pruning & Recovery is effective in removing redundant Gaussians (Exp #1$\rightarrow$3) with negligible quality degradation while preserving rendering accuracy. This proves that the proposed Global Significance, based on Gaussian attributes, accurately represents the critical aspects. By removing the high-degree SHs and transferring the knowledge to a compact representation (Exp #3 $\rightarrow$5), our method successfully demonstrates the benefits from using soft targets and extra data from view augmentation, results in negligible changes in specular reflection In practical post-processing, the vector quantization on the least important Gaussians (Exp #7), combined with the octree-based lossless compression for its geometric position (Exp #8), showcases the advantage of adopting a hybrid compression.
@@
To discover the most favorable criteria to measure the global significance of each Gaussian, we study the critical elements within each Gaussian, including its hit count with all training rays, Gaussian opacity, and the function-related volume. As shown in Tab. 3, considering all these three attributes by weighting the hit count using opacity and Gaussian volume produces the most accurate rendering quality after zero-shot pruning. A brief Gaussian Co-adaptation (last row) could restore accuracy to its level prior to pruning. We visualize the rasterized images before and after pruning, as well as the pruned Gaussians, in Fig. 6.
SH Distillation
Directly eliminating the high-degree component of SHs leads to an unacceptable performance drop from the full model (Exp #3 $\rightarrow$ 4), where we can also observe a significant loss of specular reflection on surfaces when changing viewpoints. However, by introducing knowledge distillation (Exp #4 $\rightarrow$ 5) from the full model, we can reduce its size and preserve the viewing effects. Additionally, adding pseudo-views during training (Exp #5 $\rightarrow$ 6) further demonstrates the effectiveness of guiding the training of a student model through the teacher model.
VecTree Attribute Compression
We study the effectiveness of applying quantization on Gaussian attributes, as shown in Tab. 4. We find that the Gaussian location is sensitive to a low-bit representation and thus apply lossless Octree-based compression on it. For the rest of the attributes, we first quantize them to half-precision, and applying Vector Quantization (VQ) on all attributes leads to inferior results, but this can be mitigated by using Gaussian Global Significance. The final module, VecTree Compression, performs Octree-based compression on Gaussian location, VQ on the least important SHs, and FP16 on the rest, achieving the best performance balance.
Degradation & Speed vs. Compression Ratio?
To strike a balance among model compression ratio and the reduced model size as well as the rendering speed, we study the effect of various ratios for Gaussian Pruning & Recovery, SHs Distilling, and VecTree Compression. As can be seen in the Fig 7 and Tab 5. The ratios selected represent a balanced point to optimize the trade-off between model size, FPS, and rendering quality (see Tab 5).
Figure 7: Visualization of The effect from different prune and VQ ratio.
@@
[7] [4] + 55% vector quantization 42.92MB 244 31.626 0.924 0.204
Table 5: Ablation under a mild compression ratio: With a Gaussian pruning ratio of 40% and VQ ratio of 55%, LightGaussian nearly preserves the rendering quality (SSIM decrease from 0.927 to 0.924), while achieving an $\sim$87reduction in model size and enhancing the FPS from 192 to 244.
5 Conclusion
We present LightGaussian, a novel framework that converts the heavy point-based representation into a compact format for efficient novel view synthesis. For practical use, LightGaussian explores the use of 3D Gaussians for modeling large-scale scenes and finds an effective way to identify the least important Gaussians grown by densification. In pursuing a compact Spherical Harmonics format, distillation is enhanced by synthesizing pseudo-views to generate more data for knowledge transfer. A VecTree compression post-processing effectively removes further redundancy by using Gaussian significance as an indicator. With these methods, the proposed representation reduces data redundancy by more than 15$\times$, further boosting the FPS to more than 200 FPS, with minimal rendering quality loss.
6 Acknowledgements
We would like to express our gratitude to Yueyu Hu from NYU for the invaluable discussion on our project.
A7 More Technical Details

### v5

@@ 3.3 Distilling into Compact SHs
In the uncompressed Gaussian Splat data, a considerable portion is occupied by Spherical Harmonics (SH) coefficients, necessitating (45+3) floating-point values per splat, which account for 81.3 percent of the total attribute volume. While directly diminishing the degree of SH is advantageous for reducing disk space consumption, this action results in a significant reduction in surface “shininess”, notably impacting the variance in specular reflections as the viewpoint changes.
To strike a balance between model size and scene modeling quality, we suggest transferring knowledge from well-trained high-degree SHs to their more compact, lower-degree counterparts through a distillation process. Specifically, we distill the knowledge from the teacher model, which employs full-degree SHs and remains static, to the student model, characterized by truncated lower-degree SHs. The supervision is based on the differences in the predicted pixel intensities, denoted as $\boldsymbol{C}$, between the two models:
@@
Synthesize Pseudo Views.
Specular reflection rays are reflected off the surface of an object when the viewpoints are altered. To enrich the training data, we can introduce additional viewpoints through sampling, which are capable of capturing such reflectance characteristics. Given our knowledge of the camera locations and viewing directions for all training views, we strategically sample pseudo views around each training view, adhering to a Gaussian distribution, while maintaining consistent camera viewing directions:
@@
3.4 Gaussian Attributes Vector Quantization
Vector quantization clusters the voxels into compact codebooks. While high compression rates are attainable, the task of quantizing Gaussian attributes in a point-based representation, inherently non-Euclidean, presents notable challenges. Our empirical investigations reveal that quantization applied to all elements, particularly attributes such as position, rotation, and scale, tends to incur significant accuracy losses when represented discretely, leading to a marked reduction in precision.
Vector Quantization.
We propose applying Vector Quantization (VQ) to the Spherical Harmonics (SH) coefficients, based on the assumption that a subgroup of 3d Gaussians typically exhibits a similar appearance. Fundamentally, VQ segments the Gaussians $\mathcal{G}=\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{N}\}$ (here we apply it on SH) to the K codes in the codebook $\mathcal{C}=\{\mathbf{c}_{1},\mathbf{c}_{2},\ldots,\mathbf{c}_{K}\}$, where each $\mathbf{g}_{j},\mathbf{c}_{k}\in\mathbb{R}^{d}$ and K $\ll$ N. $d$ means SH dimension. We aim to strike a balance between rendering quality loss and compression rate by leveraging the pre-computed significance score from Eq. 3. Based on this, we apply Vector Quantization (VQ) selectively on the least significant elements in the Spherical Harmonics (SHs). Specifically, we initialize $\mathcal{C}$ via K-means, iteratively sample a batch of $\mathcal{G}$, associates them to the closest codes by euclidean distance, and update each $\mathbf{c}_{k}$ via moving average rule: $\mathbf{c}_{k}=\lambda_{d}\cdot\mathbf{c}_{k}+(1-\lambda_{d})\cdot 1/\mathbf{T}_{k}\cdot\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}\mathbf{GS}_{j}\cdot\mathbf{g}_{j}$, where $\mathbf{T}_{k}=\sum_{\mathbf{g}_{j}\in\mathcal{R}(\mathbf{c}_{k})}{\mathbf{GS}_{j}}$ is the significance score (Eq. 3) which is assigned to the code vector $\mathbf{c}_{k}$, $\mathcal{R}(\mathbf{c}_{k})$ is the set of Gaussians associated to the $k$-th code. $\lambda_{d}=0.8$ represents the decay value, which is utilized to update the code vector using a moving average. We fine-tune the codebook for 5,000 iterations, while fixing the gaussian-to-codebook mapping. We disable additional clone/split operations and leverage photometric loss on the training views.
To maintain the integrity of crucial attributes such as these Spherical Harmonics with higher global significance scores, and Gaussian’s position, shape, rotation, and opacity attributes, we opt to bypass the VQ process and instead store these features directly in the float16 format.
Table 1: Quantitative Comparisons in Real-world Large-scale Scenes. We compare LightGaussian with the original 3D-GS [32], efficient voxel-based NeRFs [67, 46], voxel NeRF with vector quantization [40], and compact MLP-based NeRF [5]. Voxel-based methods all exhibit a lack of sufficient capacity for representing large-scale scenes and are not able to run at real-time speed. Mip-NeRF 360 produces the best visual quality, but it requires more than 16s to render a single image. Our method achieves a good balance among FPS, model size, and rendering quality. We re-train 3D-GS using their original code and configurations on our platform, to perform a fair comparison, noted as 3D-GS.
@@ 3.4 Gaussian Attributes Vector Quantization
Plenoxels [67] 6.79 2.1 GB 23.08 0.626 0.463 11.2 2.3 GB 21.08 0.719 0.379
INGP-Big [46] 9.43 48MB 25.59 0.699 0.331 14.4 48MB 21.92 0.745 0.305
Mip-NeRF 360 [5] 0.06 8.6MB 27.69 0.792 0.237 0.14 8.6MB 22.22 0.759 0.257
VQ-DVGO [40] 4.65 63MB 24.23 0.636 0.393 - - - - -
3D-GS [32] 134 734MB 27.21 0.815 0.214 154 411MB 23.14 0.841 0.183
3D-GS [32] 119 724MB 27.53 0.810 0.227 106 380MB 23.11 0.822 0.219
Ours 209 42MB 27.28 0.805 0.243 209 22MB 23.11 0.817 0.231
Figure 4: Visual Comparisons. We compare LightGaussian with the vanilla 3D-GS [32], presenting a residual map between prediction and ground-truth scaled from 0 to 127 to highlight the differences. We observe that LightGaussian could preserve most of the specular reflection (yellow boxes) after converting to a compact format. Additionally, a slight lightness change is noted after the conversion, as shown in the bottom white box. Note that the residual maps plotted illustrate the differences between the rendered images and the GT RGB values. Here, darker regions indicate closer alignment. For a comprehensive dynamic viewpoint comparison, please refer to our supplementary video material.
@@ 4.1 Experimental Settings
Datasets and Metrics.
We conduct comparisons using the scene-scale view synthesis dataset provided by Mip-NeRF360 [6], which comprises nine real-world large-scale scenes, including five unbounded outdoor and four indoor settings with intricate backgrounds. Furthermore, we utilize the Tanks and Temples dataset [34], a comprehensive unbounded dataset, for additional comparisons, selecting the same scenes as those used in 3D-GS [32]. Performance metrics on synthetic object-level datasets will be detailed in the supplementary materials due to constraints on space. We report metrics including the peak signal-to-noise ratio (PSNR), structural similarity (SSIM), and perceptual similarity as measured by LPIPS [70].
Compared Baselines.
We conduct comparisons with methodologies capable of modeling large-scale scenes. This includes Plenoxel [67], Mip-NeRF360 [5], and 3D-GS [32]. Additionally, we compare against efficient methods like Instant-ngp [46], which leverages a hash grid for storage efficiency, and VQ-DVGO [40]. The latter builds upon the DVGO framework [57], incorporating enhancements such as voxel pruning and vector quantization for optimized representation.
Table 2: Ablation studies on the Gaussian Pruning & Recovery, SH Compactness, and the Vector Quantization. Scene: Room. Zero-shot Gaussian pruning leads to a degradation in rendering quality (#2), but Co-adaptation can recover most of the scene details (#3). Directly eliminating high-order SH negatively affects the quality (#4), while distillation with pseudo-view helps to mitigate the gap (#5, #6). Codebook quantization further significantly reduces the required model size and bandwidth (#7).
@@
[7]    + Codebook Quant. 300.60 20MB 31.16 0.915 0.218
[8]    + VQ finetune 300.60 20MB 31.67 0.923 0.212
[9] LightGaussian (Ours) 300.60 20MB 31.67 0.923 0.212
Table 3: Ablation study of the Gaussian Pruning & Recovery, by using different Gaussian attributes for computing its global significance score. By considering only the hit count of each Gaussian from training rays, the zero-shot pruning leads to inferior performance. Incorporating the opacity and volume drives us to a better criterion. The subsequent Gaussian Co-adaptation is used to recover most of the information loss from the pruning of redundant Gaussians.
@@
Table 4: Ablation study Gaussian Attribute Vector Quantization (VQ). By quantizing all attributes to FP16, a smaller model is achieved. VQ is applied to all attributes leads to inferior modeling accuracy, but this can be mitigated by using Global Significance (GS) on the least crucial Gaussians. Other attributes (e.g., scale) are also sensitive to VQ, hence we only apply VQ on the SH. Combining VQ and significance score as indicators, LightGaussian demonstrates a good balance between size and quality.
@@
+FP16 33.56MB 31.47 0.921 0.212
   + VQ All att. 18.74MB 23.02 0.748 0.376
   + VQ All att. $\times$ GS 18.74MB 26.33 0.831 0.325
   + VQ SH. 19.89MB 30.85 0.910 0.222
   + VQ SH $\times$ GS 19.89MB 31.16 0.915 0.218
LightGaussian (w/ VQ Finetune) 19.89MB 31.67 0.923 0.212
Implementation Details.
Our framework is implemented using Pytorch and incorporates the differential Gaussian rasterization technique from 3D-GS [32]. We utilize our own trained models, ensuring they align with the performance benchmarks established in the 3D-GS study [32]. All performance are evaluated on a A6000 GPU, the same as 3D-GS. In the Global Significance Calculation phase, we assign a power value of 0.1 in Eq. 4 and proceed to fine-tune the model for 5,000 steps during the Gaussian Co-adaptation process. For SH Distillation, we downscale the 3-degree SHs to 2-degree, thereby reducing 21 elements for each Gaussian. This process is further optimized by setting $\sigma$ to 0.1 in the pseudo view synthesis stage. In the Gaussian Vector Quantization step, the codebook size is configured to 8192, selecting Spherical Harmonics (SHs) with the least $60\%$ significance score for the vector quantization (VQ ratio), to balance the trade-off between compression efficiency and fidelity.
Figure 5: Visual Comparisons for Ablation Study. We visualize the rendered RGB images and the residual map between the ground-truth image, aligned with the experiment ID as shown in Tab. 2. The final model (Exp #9) demonstrates close results to 3D-GS (Exp #1), while the Gaussian Co-adaptation, along with SH distillation, almost completely mitigates the information loss. To highlight the difference, we visualize the SSIM between the rendered images and the GT, where a whiter output in the residual signifies closer alignment with the GT, denoting better quality, and the area with greater color intensity indicates lower quality.
@@ 4.2 Experimental Results
Quantitative Results.
We evaluate the performance of various methods for novel view synthesis by summarizing their quantitative metrics in Tab. 7. This includes highly efficient voxel-based NeRFs such as Plenoxel[67] and Instant-NGP [46], the compact MLP-based Mip-NeRF360 [5], NeRF enhanced with vector quantization [40], and 3D Gaussian Splatting [32]. Specifically, within the Mip-NeRF360 dataset, NeRF-based methods leverage MLPs for a compact representation, achieving competitive accuracy. However, their slow inference speed (0.06FPS) limits practical applications. Voxel-based NeRFs improve rendering efficiency but still fall short of real-time performance. For instance, Plenoxel demands 2.1GB for storing a large-scale scene. VQ-DVGO addresses storage concerns yet faces generalization issues in complex large-scale scenes due to its reliance on pruning and vector quantization alone. Conversely, 3D-GS offers a commendable balance between rendering quality and speed but requires substantial storage for a single scene. Our method, LightGaussian, surpasses existing techniques with a rendering speed exceeding 200 FPS, courtesy of its efficient rasterization process that prunes insignificant Gaussians. Collectively, the proposed techniques significantly reduces the redundancy in 3D Gaussian model, cutting down the size from 724MB to 42MB on the Mip-NeRF360 datasets—a reduction factor of 15$\times$. LightGaussian also outperforms 3D-GS in rendering speed on the Tank & Temple datasets by nearly 2$\times$ while shrinking the storage requirement from 380MB to 22MB.
Qualitative Results.
We conducted a comparative analysis of the rendering outcomes between 3D-GS and our LightGaussian model, focusing particularly on the intricate details and background regions, as demonstrated in Fig. 4. It is evident that both 3D-GS and LightGaussian maintain comparable visual quality, even in scenes with thin structures (refer to the first and third rows).
@@ 4.3 Ablation Studies
We conducted ablation studies on our method’s components to individually assess the impact of the proposed modules. Our findings reveal that Gaussian Pruning and Recovery is effective in eliminating redundant Gaussians, ensuring that only significant ones contribute to the scene representation. SH Distillation proves to be effective in reducing the degree of Spherical Harmonics, which simplifies the lighting model without compromising too much quality. Additionally, Vector Quantization demonstrates its efficiency in compressing the feature space, resulting in a more streamlined and efficient representation. When combined, these modules significantly enhance the overall efficacy and performance of our framework.
Overall Analysis.
We report the performance of the investigation on the proposed modules in Tab. 2 and Fig. 5. We verify that our design of Gaussian Pruning & Recovery is effective in removing redundant Gaussians (Exp #1$\rightarrow$3) with negligible quality degradation while preserving rendering accuracy. This proves that the proposed global significance, based on principle of splatting, accurately represents the critical aspects of each Gaussian. By removing the high-degree SHs and transferring the knowledge to a compact representation (Exp #3 $\rightarrow$5), our method successfully demonstrates the benefits from using soft targets and extra data from view augmentation, results in negligible changes in specular reflection. In practical post-processing, the vector quantization on the least important Gaussians (Exp #7), showcases the advantage of adopting VQ to further reduce model size.
@@
To discover the most favorable criteria to measure the global significance of each Gaussian, we evaluated key characteristics within each Gaussian: the Gaussian/rays interaction count, Gaussian opacity, and the volume related to its function. Tab. 3 illustrates that integrating these three factors—by weighting the hit count with opacity and Gaussian volume—yields the highest rendering quality post zero-shot pruning. A quick Gaussian Co-adaptation (the last row) which optimizes the remaining significant Gaussians, can effectively recover the rendering accuracy to its pre-pruning level. The visual comparison of rasterized images, before and after pruning, alongside the depiction of pruned Gaussians, is presented in Fig. 6.
SH Distillation & Vector Quantization.
The direct removal of high-degree components in Spherical Harmonics (SHs) results in a substantial performance decline compared to the full model (Exp #3 $\rightarrow$ 4), manifesting notably in the loss of specular reflections on surfaces with varying viewpoints. However, the integration of knowledge distillation (Exp #4 $\rightarrow$ 5) from the full model enables us to decrease its size while maintaining the essential viewing effects. Furthermore, the incorporation of pseudo-views during the training process (Exp #5 $\rightarrow$ 6) validate the efficacy of using the simulated views to enhance the learning of specular reflections. Applying Vector Quantization (VQ) on all attributes leads to inferior results (See Tab. 4), but this can be mitigated by using Gaussian Global Significance.
Table 5: Ablation under different SH ompactness: With a Gaussian pruning ratio of 66% and reduce SH to 2-degree, LightGaussian nearly preserves the rendering quality (SSIM decrease from 0.927 to 0.926). Reducing the SH to 1-degree lead to the SSIM decreases to 0.923.
@@
Figure 7: Visualization of the rendering performance vs. different prune and VQ ratios. The quantize ratio is the compression ratio at vector quantization.
Degradation & Speed vs. Compression Ratio
We investigate the interplay between rendering quality and speed across various model compression ratios. Specifically, we adjust parameters such as Gaussian Pruning & Recovery, SH Distilling, and VQ. As depicted in Fig 7, we note a marked decline in rendering quality when the pruning ratio reaches 70%, with a more rapid deterioration as the VQ ratio approaches 65%. In Tab 5, we observe that 2-degree SH compactness reduces the SSIM from 0.927 to 0.926, while 1-degree SH compactness further reduces it to 0.923.
5 More Technical Details
Algorithm Explanation.